Use algebraic procedures to find the exact solution(s) of the equation. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.)

e^(x + 3) = 22
x=?

Answer by ankor@dixie-net.com(22740)

e^(x + 3) = 22
Find the nat log of each side
ln(e^x+3)) = ln(22)
log equiv of exponents
(x+3)ln(e) = ln(22)
Find the log, ln of e is 1 so we have
x + 3 = 3.09104
x = 3.09104 - 3
x = .09104
:
:
Check this on you calc
enter e^(.09104+3) results: 21.99999 ~ 22